OpenStudy (anonymous):

identify the center and radius of the circle (x-5)^2+(y+1)^2=4

OpenStudy (danjs):

c(5,-1) r = 2

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

r is radius so for your equation r^2=4, so r=2

OpenStudy (jhannybean):

@DanJS provide explanations rather than giving answers.

OpenStudy (anonymous):

center is (h, k), so it will be (5, -1)

OpenStudy (anonymous):

center (5, -1) radius=2

OpenStudy (danjs):

@Jhannybean sorry, beg my pard

OpenStudy (anonymous):